# Neural Algorithmic Reasoning with Multiple Correct Solutions

### v3

@@ 3.2 Results for Depth-First Search
True (n = 5) $1.00\pm 0.00$ $1.00\pm 0.00$ $1.00\pm 0.00$ $00.00\pm 0$ 0.00
Model (n = 5) $1.00\pm 0.00$ $1.00\pm 0.00$ $1.00\pm 0.00$ $00.00\pm 0$ 0.00
True (n = 64) $1.00\pm 0.00$ $1.00\pm 0.00$ $1.00\pm 0.00$ $00.00\pm 0$ 0.00
Model (n = 64) $0.50\pm 0.28$ $0.56\pm 0.18$ $0.38\pm 0.19$ $00.00\pm 0$ 0.00
@@
True (n = 5) $0.80\pm 0.10$ $0.90\pm 0.07$ $0.34\pm 0.18$ $00.00\pm 0$ 0.00
Model (n = 5) $0.72\pm 0.28$ $0.54\pm 0.11$ $0.14\pm 0.15$ $00.00\pm 0$ 0.00
True (n = 64) $0.10\pm 0.07$ $00.00\pm 0$ 0.00 $0.04\pm 0.05$ $00.00\pm 0$ 0.00
Model (n = 64) $0.14\pm 0.21$ $0.04\pm 0.05$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00
Conclusion
We outline a two-part approach to NAR with multiple solutions: we train models to predict distributions of solutions and stochastically extract solutions from those distributions. We demonstrate feasibility, given a good distribution, but also that NNs do not easily learn to predict good distributions. We attribute this difficulty to a number of factors. First, NAR is difficult: previous work has shown that NNs only achieve up to 40% micro-F1 on deterministic DFS without hint pointer reversal, which improves it to nearly 100% [4]. Second, training a model to fit a distribution of solutions does not directly optimize for correct solutions. Solutions must still be extracted from the distribution via stochastic methods, to which the loss metric does not extend (Figure 7). Improving methods for extracting solutions remains an important algorithmic task. Still, with the promise of sampling multiple solutions from a distribution demonstrated, we believe our modifications of the CLRS benchmark act as proof-of-concept for further work. Our method is applicable to other CLRS-30 algorithms such as Minimum Spanning Tree, Bipartite Matching, and Task Scheduling. Having multiple solutions adds essential flexibility to applications of NAR such as cyberphysical systems and might aid in explaining NAR models.
@@ Appendix A Methodology
We discuss our methods for randomising the deterministic (canonical) versions of Depth-First Search and the Bellman-Ford algorithm in order to use the resulting distributions over solutions as inputs to the CLRS-30 baseline. We also present our methods of extracting candidate solutions from model outputs and give approaches towards verifying them as correct solutions of the respective deterministic algorithms our sampling methods are aiming to emulate.
@@
Figure 2: Parent Tree encoding of a single DFS solution. The start node, 0, is its own parent, represented by the value 0 at index 0. The value 2 at index 1 indicates that vertex 2 is the parent of vertex 1. Similarly, the value 0 at index 2 indicates that vertex 0 is the parent of vertex 2.
@@
Figure 3: Parent Probability Distribution encoding of multiple DFS solutions. The value 0.5 in row $1$ column $2$ indicates that vertex 2 is a parent of vertex 1 in 50% of correct parent trees. We call the parent probabilities obtained by repeatedly running an algorithm the true parent distribution $\mathcal{P}$. The model outputs, which we aim to be as close as possible to $\mathcal{P}$, are denoted as $\tilde{\mathcal{P}}$.
@@
Figure 4: DFS tree extraction from a parent probability matrix: $\tilde{\mathcal{P}}$ or $\mathcal{P}$. The parent tree [0,2,2] is incorrect because there is no solution in which 2 is its own parent. [0,0,0] is incorrect, despite there being solutions in which 0 is the parent of each vertex, because there is no solution in which 0 is the parent of every vertex: if 0 is the parent of 1, 1 must be the parent of 2.
A.1 DFS
@@
We consider a generalization of the deterministic DFS setting implemented in the CLRS benchmark [2]. In our setting, edge-exploration is random: from a given vertex, any of the outgoing edges may be explored first. Still, we require that once DFS backtracks at the root of a connected component, it restarts its search at the next vertex in a linear ordering of the vertices. We think the generalization is natural, modelling a setting where one conducts DFS according to a chosen sequence of starting points, without assumptions about which edges to pursue first.
@@
Figure 5: A graph with multiple correct DFS solutions. Graph edges are drawn in black. The red tree would be the result if vertices were explored in alphabetical order, while the blue tree is one of multiple other posible valid DFS trees. Note also how the DFS-tree is encoded as a predecessor array (with the entry at position $i$ is the predecessor of the $i^{th}$ vertex in alphabetical order, which we henceforth call $\pi$.
A.1.1 Randomness in training
We generate a ground-truth parent distribution $\mathcal{P}$ for each graph in the test set, and train the model to minimize the KL-divergence between its predicted distribution $\tilde{\mathcal{P}}$ and $\mathcal{P}$. We obtain $\mathcal{P}$ by combining 20 solutions generated by a DFS algorithm that randomly tiebreaks between outgoing edges. Refer to Appendix B for a justification of using 20 solutions. We compute the fraction of solutions in which node $s$ is the parent of node $t$, which we say is the parent probability, as shown in Figure 3. Algorithm 1 gives detailed description of our particular DFS algorithm, with the line that randomizes edge-exploration highlighted in red.
@@ A.1.1 Randomness in training
3:Initialize color $\leftarrow[0]*|V|$triangleright list of |V|-many 0s
@@ A.1.2 Randomness in sampling
$u=\underset{v\in V(G)}{\text{max}}\tilde{\mathcal{P}}_{v,u},$ (1)
where $\tilde{\mathcal{P}}_{v,u}$ is the probability of an edge $(v,u)$ according to $\tilde{\mathcal{P}}$. Argmax sampling is deterministic and can return only one solution. We include it in our discussion as a good baseline which corresponds to the internal mechanism of the CLRS-30 benchmark [2]. To extract multiple solutions for DFS, we introduce stochasticity according to Upwards Sampling, which is described in Algorithm 2.
@@
6:     leaf $\leftarrow leaves[0]$triangleright Take the "leafiest" vertex
@@
In Upwards Sampling (Algorithm 2), we first determine how likely each node is to be a parent node of any other node. We do so by summing the probabilities of each vertex $u\in V$ to be the parent of each vertex $v\in V$ and then sorting this list in ascending order, as done in Line 4. In each iteration of the outer loop (Line 5), we begin by taking the "leafiest" vertex (the vertex with the lowest probability of being a parent) and sampling its parent according to $\tilde{\mathcal{P}}_{leaf}$ (Line 7). After having done so, we remove the leaf from future consideration as a leaf and as a parent (as a leaf by definition will not be the parent of another node). Next, we sample the parent of $\pi[leaf]$ until we encounter a node that already has a parent (Line 5). This ensures we do not overwrite previously sampled parents. We have chosen this sampling method as it works similarly to DFS while respecting the inherent stochasticity resulting from returning $\tilde{\mathcal{P}}$ rather than one unambigous solution, as has been done in [2] and [3]. During testing, we encountered the situation that the removal of leaves from consideration as parents leads to no potential parents being available for a node due to having been removed. To address this problem, we give a node a random parent in this case. We further implement a method without this removal, which we call alt Upwards Sampling. While less faithful to graph properties, it addresses the limitations caused by the difficulty of computing the "leafiness" of each node while operating closer to DFS. It can be seen in Section 3.2 that this method outperforms Algorithm 2.
A.1.3 Checking Validity
Reconstructing multiple solutions from a matrix of probabilities introduces the possibility of error. We test the solutions we extract from our models against the following necessary conditions for correctness.
We say purported parent tree $\pi:=(v_{0},v_{1},...,v_{|V|-1})$ is DFS-valid for a graph $G=(V,E)$ with linearly ordered vertices $V$ only if the following 5 conditions hold:
  • The start node is correct:
    The least element of the linear ordering on Vertices is its own parent: $\pi[0]:=v_{0}=0=V[0]$.
  • The edges are correct:
    All edges suggested by the parent tree are in fact edges in the graph:
    $\forall t\in V,t==\pi[t]\lor E(\pi[t],t)$.
  • There is no cycle bigger than a self-loop in the parent tree:
    $\neg\exists C\subseteq V:\forall v\in C,\exists s,t\in C,s\neq v,t\neq v:E(s,v)\land E(v,t).$
  • Each vertex which is its own parent is not reachable by a lower-index node:
    $\forall t\in V,\pi[t]==t\implies\forall s\in V,s<t,\neg path(s,t)$.
  • Each parent whose child is not itself is reachable from the lowest index node for which its child is reachable: $\forall t\in V,\pi[t]\neq t\implies\exists l\in V:path(l,\pi[t])\land\forall s\in V,path(s,t)\implies l\leq s$111path(x,x) is true
No sufficient conditions to determine whether, given a tree $T$ and a graph $G$, $T$ is a valid DFS-tree of $G$ (for some tie-breaking rule) are known to the authors in the graph theory literature.
A.2 Bellman-Ford
@@ A.2.1 Randomness in training
Figure 6: A graph with multiple possible minimum-cost $(source,v)$-paths. As previously, the black arrows represent the edges, with the red and blue edges showing different minimum-cost paths corresponding to their respective predecessor arrays.
A.2.2 Randomness in sampling
We use an algorithm we call BF Beamsearch to sample predecessor arrays $\pi$ representing the minimum-cost $(s,v)$-paths for each $v\in V(G)$. This algorithm closely resembles standard Beamsearch, but it differs as it samples according to the distribution given by the model. For each $v\in V(G)$, BF Beamsearch considers paths from the source to $v$ of length at most $n$ by first sampling candidate predecessors of $v$ and then recursively sampling predecessors of predecessors until the path has length $n$. At each stage, only the 3 lowest-cost paths are kept and used for the next sampling. As soon as the maximum recursion depth is reached, the predecessor that is at the end of the lowest-cost path from the source is chosen as $\pi[v]$. This mimics the operation of the deterministic Bellman-Ford algorithm while respecting randomness and is thus a suitable, albeit involved, way of sampling $\pi$. To sample predecessor arrays more easily from $\mathcal{P}$ and $\tilde{\mathcal{P}}$, we use a greedy randomised algorithm. This algorithm samples a set of parents for each vertex $v$ except the source node (which we assume to be its own parent) according to $\tilde{\mathcal{P}}_{v}$. In the next step, it checks whether any of the sampled parents are plausible (i.e., if there exists an edge $(p,v)\in E(G)$ for the parent $p$). It resamples if there are no plausible parents. If any of the sampled parents are plausible, or if the maximum number of resamplings is reached, it chooses the plausible parent $p$ with the lowest weight $w_{p,v}$ as the predecessor $\pi[v]$. Furthermore, we also use the methods of deterministically choosing the likeliest parent of $v$ according to $\tilde{\mathcal{P}}_{v}$ and randomly uniformly choosing a parent as was done with DFS in order to obtain some baseline results.
A.2.3 Checking Validity

### v4

@@ 3.2. Results for Depth-First Search
True (n=5) $1.00\pm 0.0$ $1.00\pm 0.0$ $1.00\pm 0.0$ $0.00\pm 0$ 0.0
Model (n=5) $1.00\pm 0.0$ $1.00\pm 0.0$ $1.00\pm 0.0$ $0.00\pm 0$ 0.0
True (n=16) $1.00\pm 0.0$ $1.00\pm 0.0$ $1.00\pm 0.0$ $0.00\pm 0$ 0.0
Model (n=16) $0.98\pm 0.0$ $0.91\pm 0.0$ $0.92\pm 0.0$ $0.00\pm 0$ 0.0
True (n=64) $1.00\pm 0.0$ $1.00\pm 0.0$ $1.00\pm 0.0$ $0.00\pm 0$ 0.0
Model (n=64) $0.50\pm 0.3$ $0.56\pm 0.2$ $0.38\pm 0.2$ $0.00\pm 0$ 0.0
@@
True (n=5) $0.80\pm 0.1$ $0.90\pm 0.1$ $0.34\pm 0.2$ $0.00\pm 0$ 0.0
Model (n=5) $0.61\pm 0.1$ $0.50\pm 0.1$ $0.16\pm 0.0$ $0.00\pm 0$ 0.0
True (n=16) $0.00\pm 0$ 0.0 $0.00\pm 0$ 0.0 $0.08\pm 0.0$ $0.00\pm 0$ 0.0
Model (n=16) $0.00\pm 0$ 0.0 $0.00\pm 0$ 0.0 $0.00\pm 0$ 0.0 $0.00\pm 0$ 0.0
True (n=64) $0.00\pm 0$ 0.0 $0.00\pm 0$ 0.0 $0.01\pm 0.0$ $0.00\pm 0$ 0.0
Model (n=64) $0.00\pm 0$ 0.0 $0.00\pm 0$ 0.0 $0.00\pm 0$ 0.0 $0.00\pm 0$ 0.0
3.3. Results for Permuting
Permuting node indices either fails to produce new solutions (each time predicting a permutation of the same original solution), or else incurs significant accuracy penalties (Table 3). For DFS with ordered restarts, some of the decline is explained because a solution that was correct originally may become incorrect after a permutation. However, for the Bellman-Ford algorithm, the solutions generated by the model remain correct after permuting the vertex indices. Even so, the neural algorithmic reasoner does not appear to be capable of generating any solutions that are correct and non-isomorphic to the single correct solution generated by the model.
Table 3. Mean accuracies and proportions of distinct solutions (0 to 1) when permuting inputs 5 times, with standard deviations over 5 runs.
Accuracy
DFS Bellman-Ford
(n=5) $0.10\pm 0.03$ $0.30\pm 0.07$
(n=16) $0.01\pm 0.02$ $00.00\pm 0$ 0.00
(n=64) $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00
Variety
DFS BF
(n=5) $0.90\pm 0.09$ $0.62\pm 0.08$
(n=16) $1.00\pm 0.00$ $0.99\pm 0.01$
(n=64) $1.00\pm 0.00$ $1.00\pm 0.00$
4. Conclusion
We outline a two-part approach to NAR with multiple solutions: we train models to predict distributions of solutions and stochastically extract solutions from those distributions. In the case of using a NAR to find multiple single-source shortest path solutions in a weighted graph we demonstrate that this is feasible, particularly on graphs with sizes that lie within the range of sizes we train our models on. For DFS, we see that this task is significantly more challenging, as indicated by prior NAR research. We attribute this result to a number of factors. First, previous work indicates that emulating DFS with NAR is a difficult task. Second, training a model to fit a distribution of solutions does not directly optimize for correct solutions. Solutions must still be extracted from the distribution via stochastic methods, to which the loss metric does not extend (Figure 12). Improving methods for extracting solutions remains an important algorithmic task. Still, with the promise of sampling multiple solutions from a distribution demonstrated, we believe our modifications of the CLRS benchmark act as proof-of-concept for further work. Our method would be applicable to other CLRS benchmark algorithms such as Minimum Spanning Tree, Bipartite Matching, and Task Scheduling. Having multiple solutions adds essential flexibility to applications of NAR such as cyberphysical systems and might aid in explaining NAR models.
@@ Appendix A Methodology
We discuss our methods for randomising the deterministic (canonical) versions of Depth-First Search and the Bellman-Ford algorithm in order to use the resulting distributions over solutions as inputs to the CLRS baseline. We also present our methods of extracting candidate solutions from model outputs and give approaches towards verifying them as correct solutions of the respective deterministic algorithms our sampling methods are aiming to emulate.
@@
Figure 7. Parent Tree encoding of a single DFS solution. The start node, 0, is its own parent, represented by the value 0 at index 0. The value 2 at index 1 indicates that vertex 2 is the parent of vertex 1. Similarly, the value 0 at index 2 indicates that vertex 0 is the parent of vertex 2.
@@
Figure 8. Parent Probability Distribution encoding of multiple DFS solutions. The value 0.5 in row $1$ column $2$ indicates that vertex 2 is a parent of vertex 1 in 50% of correct parent trees. We call the parent probabilities obtained by repeatedly running an algorithm the true parent distribution $\mathcal{P}$. The model outputs, which we aim to be as close as possible to $\mathcal{P}$, are denoted as $\tilde{\mathcal{P}}$.
@@
Figure 9. DFS tree extraction from a parent probability matrix: $\tilde{\mathcal{P}}$ or $\mathcal{P}$. The parent tree [0,2,2] is incorrect because there is no solution in which 2 is its own parent. [0,0,0] is incorrect, despite there being solutions in which 0 is the parent of each vertex, because there is no solution in which 0 is the parent of every vertex: if 0 is the parent of 1, 1 must be the parent of 2.
A.1. DFS
@@
We consider a generalization of the deterministic DFS setting implemented in the CLRS benchmark (Veličković et al., 2022). In our setting, edge-exploration is random: from a given vertex, any of the outgoing edges may be explored first. Still, we require that once DFS backtracks at the root of a connected component, it restarts its search at the next vertex in a linear ordering of the vertices. We think the generalization is natural, modelling a setting where one conducts DFS according to a chosen sequence of starting points, without assumptions about which edges to pursue first.
@@
Figure 10. A graph with multiple correct DFS solutions. Graph edges are drawn in black. The red tree would be the result if vertices were explored in alphabetical order, while the blue tree is one of multiple other posible valid DFS trees. Note also how the DFS-tree is encoded as a predecessor array (with the entry at position $i$ is the predecessor of the $i^{th}$ vertex in alphabetical order, which we henceforth call $\pi$.
A.1.1. Randomness in training
Algorithm 1 gives detailed description of the adapted randomized DFS algorithm used in this paper, with the line that randomizes edge-exploration highlighted in red. After running Algorithm 1 20 times, we compute the fraction of solutions in which node $s$ is the parent of node $t$, which we say is the parent probability, as shown in Figure 3.
@@ A.1.1. Randomness in training
3:Initialize color $\leftarrow[0]*|V|$
@@ A.1.2. Randomness in sampling
(1) $u=\underset{v\in V(G)}{\text{max}}\tilde{\mathcal{P}}_{v,u},$
where $\tilde{\mathcal{P}}_{v,u}$ is the probability of an edge $(v,u)$ according to $\tilde{\mathcal{P}}$. Argmax sampling is deterministic and can return only one solution. We include it in our discussion as a good baseline which corresponds to the internal mechanism of the CLRS-30 benchmark (Veličković et al., 2022). To extract multiple solutions for DFS, we introduce stochasticity according to Upwards Sampling, which is described in Algorithm 2.
@@
6:     leaf $\leftarrow leaves[0]$triangleright Take the vertex most likely to be a leaf
@@
In Upwards Sampling (Algorithm 2), we first determine how likely each node is to be a parent node of any other node. We do so by summing the probabilities of each vertex $u\in V$ to be the parent of each vertex $v\in V$ and then sorting this list in ascending order, as done in Line 4. In each iteration of the outer loop (Line 5), we begin by taking the vertex with the lowest probability of being a parent and sampling its parent according to $\tilde{\mathcal{P}}_{leaf}$ (Line 7). After having done so, we remove the leaf from future consideration as a leaf and as a parent (as a leaf by definition will not be the parent of another node). Next, we sample the parent of $\pi[leaf]$ until we encounter a node that already has a parent (Line 5). This ensures we do not overwrite previously sampled parents. We have chosen this sampling method as it works similarly to DFS while respecting the inherent stochasticity resulting from returning $\tilde{\mathcal{P}}$ rather than one unambigous solution, as has been done in (Veličković et al., 2022) and (Ibarz et al., 2022). During testing, we encountered the situation that the removal of leaves from consideration as parents leads to no potential parents being available for a node due to having been removed. To address this problem, we give a node a random parent in this case. We further implement a method without this removal, which we call AltUpwards Sampling. While less faithful to graph properties, it addresses the limitations caused by the difficulty of computing the likelihood of being a leaf for each node while operating closer to DFS. It can be seen in Section 3.2 that this method outperforms Algorithm 2.
A.1.3. Checking Validity
We verify whether a given forest $F$ is the result of some depth-first traversal of the graph with ordered restarts (RDFSO). By RDFSO, we mean DFS with randomised tiebreaking and ordered restarts. Randomised tiebreaking means that given two successors $v_{1},v_{2}$ of a node $v\in G$, the algorithm chooses randomly from which node to continue the depth-first traversal. Ordered restarts means that when all descendants of a given root have been visited by the search, the algorithm continues DFS from the lowest-indexed unvisited node. Since we are not aware of a solution to this problem in the literature, we define it as the randomised DFS recognition problem (RDFSR) and present an algorithm to solve it.
Theorem A.1.
Algorithm 3 accepts if and only if a forest $F$ is the output of some run of randomised DFS with ordered restarts (RDFSO) on $G$.
Idea 0.
The algorithm attempts to build a valid DFS traversal of $G$, using $F$ to inform the traversal order. We derive the traversal order by comparing the descendants of each node in the graph $G$ and forest $F$. The key fact is that any depth-first traversal from node $v$ visits all unvisited nodes reachable from $v$ before backtracking to $v$. As a result, all unvisited nodes reachable from $v$ in graph $G$ are still reachable from $v$ in DFS-forest $F$. We use this to rebuild a DFS traversal for $F$ one vertex at a time, visiting next whichever vertex has the same unvisited descendants in $G$ that it does in $F$.
We present Algorithm 3, which solves DFS verification by rebuilding $F$ by traversing $G$. Algorithm 3 rebuilds $F$ one vertex at a time, removing each vertex after considering it. Algorithm 3 discovers the next vertex in traversal by calling Algorithm 4.
Algorithm 4, the subroutine called by our main algorithm, finds the next vertex after $v$ in the traversal order by looking for the adjacent vertex to $v$ in $F$ that discovers all its descendants in $G$. If our most recently visited node $m$ has two children $v_{1},v_{2}$ in $F$, Algorithm 4 determines whether traversal can continue first to $v_{1}$ or $v_{2}$ and still produce $F$.
Line 2 checks whether the vertex $v$ is indeed a valid next vertex for the traversal. We write $D_{G}(v)$ to mean the descendants of node $v$ in graph $G$. Line 2 checks whether $D_{G}(v)=D_{F}(v)$.
The loop in Lines 7-18 of Algorithm 4 finds the first vertex $v_{1}$ adjacent to $v$ for which $D_{G}(v_{1})=D_{F}(v_{1})$ and recursively repeats this procedure to retrace the run of RDFSO until it reaches the maximum depth in $G$, removing the vertices discovered during the recursive calls in $G$ and $F$, thus rendering them unreachable in $G$ from other vertices. It then continues the loop in Lines 7-18 with the remaining descendants of $v$.
Say there is a vertex $v_{2}$ adjacent to $v$ in $F$ from which the depth-first traversal has continued after having terminated its traversal from $v_{1}$. After the vertices discovered by $v_{1}$ have been removed from $G$, the sets of vertices reachable from $v_{2}$ in $F$ and $G$ must now be the same. This is the case because if $D_{G}(v_{1})\cap D_{G}(v_{2})\neq\emptyset$ before the traversal, then all vertices in $D_{G}(v_{1})\cap D_{G}(v_{2})$ have been removed during the call from $v_{1}$ (or during an earlier call) and are now unreachable from $v_{2}$ in $G$. Our algorithm now picks $v_{2}$ as the next vertex from which to continue the traversal. Therefore, if Algorithm 4 has not rejected previously, then the set of vertices reachable from $v_{2}$ in $F$ and $G$ at a given point could have only been discovered in a traversal from $v_{2}$. We thus continue the recursive traversal from $v_{2}$, et cetera.
If the algorithm encounters a situation where none of the remaining descendants of a given vertex can reach the same set of reachable vertices in $F$ and $G$, the algorithm rejects. This is equivalent to the idea that there is no vertex the traversal can continue towards and still result in $F$. Therefore, the algorithm rejects if there is ever no vertex that discovers all its descendants in $G$ during the run of RDFSO and accepts only if such a vertex exists at every depth of the recursion.
Algorithm 3 RDFSO Verifier
1:Input: $G,F$
2:for each vertex $i$ in $G$ do
3:     if not DVC($G$, $F$, $i$then
4:         return False
5:     end if
6:end for
7:return True
Algorithm 4 Descendant Validity Check (DVC)
1:Input: $G,F,\text{vertex }v$
2:if $D_{G}(v)\neq D_{F}(v)$ then
3:     return False
4:end if
5:$descendants\leftarrow$ descendants of $v$ in $F$
6:Delete $v$ from $G$ and $F$
7:for each $descendant$ in $descendants$ do
8:     Compute $pd[descendant]$triangleright compute the possible descendants in $G$
9:     Compute $ad[descendant]$triangleright compute the actual descendants in $F$
10:     if $pd[descendant]=ad[descendant]$ then $\triangleright$ next vertex in traversal found
11:         if not DVC($G$, $F$, $descendant$then
12:              return False
13:         end if
14:         restart line 7, with updated $descendants$triangleright vertices may have been deleted
15:     else
16:         Check next descendant
17:     end if
18:end for
19:return True if $descendants$ is empty. Else return False.
A.2. Bellman-Ford
@@ A.2.1. Randomness in training
Figure 11. A graph with multiple possible minimum-cost $(source,v)$-paths. As previously, the black arrows represent the edges, with the red and blue edges showing different minimum-cost paths corresponding to their respective predecessor arrays.
A.2.2. Randomness in sampling
For each $v\in V(G)$, BF Beamsearch considers paths from the source to $v$ of length at most $n$ by first sampling candidate predecessors of $v$ and then recursively sampling predecessors of predecessors until the path has length $n$. At each stage, only the 3 lowest-cost paths are kept and used for the next sampling. As soon as the maximum recursion depth is reached, the predecessor that is at the end of the lowest-cost path from the source is chosen as $\pi[v]$. This mimics the operation of the deterministic Bellman-Ford algorithm while respecting randomness and is thus a suitable, albeit involved, way of sampling $\pi$. To sample predecessor arrays more easily from $\mathcal{P}$ and $\tilde{\mathcal{P}}$, we use a greedy randomised algorithm. This algorithm samples a set of parents for each vertex $v$ except the source node (which we assume to be its own parent) according to $\tilde{\mathcal{P}}_{v}$. In the next step, it checks whether any of the sampled parents are plausible (i.e., if there exists an edge $(p,v)\in E(G)$ for the parent $p$). It resamples if there are no plausible parents. If any of the sampled parents are plausible, or if the maximum number of resamplings is reached, it chooses the plausible parent $p$ with the lowest weight $w_{p,v}$ as the predecessor $\pi[v]$. Furthermore, we also use the methods of deterministically choosing the likeliest parent of $v$ according to $\tilde{\mathcal{P}}_{v}$ and randomly uniformly choosing a parent as was done with DFS in order to obtain some baseline results.
A.2.3. Checking Validity